# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v1

@@ Abstract.
High-Level Synthesis enables the rapid prototyping of hardware accelerators, by combining a high-level description of the functional behavior of a kernel with a set of micro-architecture optimizations as inputs. Such pragmas may describe the pipelining and replication of units, or even higher level transformations for HLS such as automatic data caching using the AMD/Xilinx Merlin compiler. Selecting the best combination of pragmas, even within a restricted set, remains particularly challenging and the typical state-of-practice uses design-space exploration to navigate this space. But due to the highly irregular performance distribution of pragma configurations, typical DSE approaches are either extremely time consuming, or operating on a severely restricted search space.
In this work we propose a framework to automatically insert HLS pragmas in regular loop-based programs, supporting pipelining, unit replication (coarse- and fine-grain), and data caching. We develop a simple analytical performance and resource model as a function of the input program properties and pragmas inserted. We prove this model provides a lower bound on the actual performance for any possible configuration. We then encode this model as a Non-Linear Program, by making the pragma configuration as unknowns of the system, which is computed optimally by solving this NLP. This approach can also be used during DSE, to quickly prune points with a (possibly partial) pragma configuration, employing this latency lower bound property. We extensively evaluate our end-to-end fully implemented system, showing it can effectively manipulate spaces of billions of designs in seconds to minutes for the kernels evaluated.
@@
copyright: noneconference: Proceedings of the 2024 ACM/SIGDA International Symposium on Field Programmable Gate Arrays; March 3 –- 5, 2024 ; Monterey, CAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
High-level synthesis (HLS) (Cong et al., 2011; Zhang et al., 2008) compilers (Xilinx, 2023b; Intel, 2023; Siemens, 2023; Microchip, 2023) and source-to-source compiler for HLS (Xilinx, 2023a; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020) can reduce development time while delivering a good performance for the designs. However, achieving a satisfactory Quality of Results (QoR) often requires design-space exploration (DSE). This is because the design space, including which pragmas to insert and where, can not only contain millions of points, but typically does not present characteristics suitable for fast analytical exploration, such as convexity and regularity. Although the existing DSE methods (Sohrabizadeh et al., 2021, 2022; Wu et al., 2023) can find designs with a good QoR, it comes at a high computation cost: for example hundreds, of designs may be concretely instantiated using HLS to compute its estimated QoR during exploration (Sohrabizadeh et al., 2021).
@@
Our main objective is to provide a system to automatically insert a set of hardware pragmas for HLS that delivers a good QoR but significantly reduces the search time needed to obtain the final design. To address this challenge, we propose NLP-DSE, a framework built on top of the AMD/Xilinx Merlin source-to-source compiler (Xilinx, 2023a). This framework automatically inserts pragmas for unrolling/parallelization, pipelining, tiling and data caching for linear programs prior to HLS. These Merlin pragmas can also be inserted using a DSE approach, such as AutoDSE (Sohrabizadeh et al., 2021), which we utilize as reference for our evaluations. However, in contrast to AutoDSE (Sohrabizadeh et al., 2021), we develop a hybrid analytical approach combined with a lightweight DSE, reducing the number of designs explored while preserving, and often even improving, the final QoR of designs produced.
@@
To this end, we create a novel Non-Linear Programming approach to automatically insert pragmas in an existing program. We develop an analytical model combining latency and resources, targeting regular loop-based kernels (PolyBench, [n. d.]), that is parameterized by the pragma configuration. We can then designate the pragma configuration as the unknowns of this model, solving it by NLP to obtain the set of pragmas that minimizes latency. An important design principle of our approach is to ensure that the latency computed is a lower bound on the achievable latency for a given pragma configuration. This enables efficient pruning during DSE: any design predicted to have a latency lower bound higher than the best latency obtained through exploration so far is necessarily slower and does not need exploration. To overcome the fact that optimizing compilers, and the overall HLS toolchain underneath, may not apply optimizations as expected (e.g., due to insufficient resources, or limitations of the compiler’s implementation), we develop a lightweight NLP-based DSE approach, exploring parts of the design space with different types and amounts of hardware parallelisms and array partitioning factors. We make the following contributions:
@@
  • We produce an analytical performance and resource model specifically built for AMD/Xilinx Vitis and Merlin compilers, which is amenable to optimization via non-linear programming.
@@
  • We prove our model is a lower bound on the final latency of the design, under reasonable hypothesis. This enables fast pruning of the design space, ensuring pruned designs have higher latency.
@@
  • We develop a NLP-based DSE approach exploiting this model, targeting regular loop-based kernels, which can significantly outperform DSE-based search approaches such as AutoDSE, delivering equal or better QoR in a significantly less search time.
@@
  • We implement NLP-DSE as an end-to-end fully automated system and use it to conduct extensive evaluation on 47 benchmarks including kernels from linear algebra, image processing, physics simulation, graph analytics, datamining, etc. (PolyBench, [n. d.]). Our results show the ability of our approach to find in most cases a better QoR than AutoDSE, in significantly less time.
@@
The paper is organized as follows. Section 2 motivates our approach and solution proposed. Section 3 presents our analytical performance and resource model. Section 4 delves into proving it is a lower bound on the final QoR. In Section 5, we introduce a non-linear formulation based on this model to automatically find pragma configurations by NLP optimization. Section 6 presents our lightweight DSE approach. Finally, Sections 7 and 9 are devoted to evaluating our method validating the effectiveness of our approach and presenting related work, before concluding.
@@ 2. Background and Motivation
The main objective of this work is to automatically optimize FPGA designs using HLS (Cong et al., 2011; Zhang et al., 2008). HLS has made FPGA usage more accessible, and many projects are looking to further democratize this field by automating optimizations (Sohrabizadeh et al., 2021; Lai et al., 2019; Huang et al., 2021; Ye et al., 2022; Ferretti et al., 2018) such as the AMD/Xilinx Merlin Compiler ((Cong et al., 2016b, a; Xilinx, 2023a)).
In particular, Merlin was developed to improve the performance and reduce the development time of HLS-based designs. To achieve this goal, it automatically generates data transfers between off-chip and on-chip memory and inserts important Vitis pragmas, such as for array partitioning. Merlin includes hardware directives pragmas such as ACCEL parallel <factor=x>, which creates x parallel instances, and Merlin restructures the code accordingly if the loop nest has more than x iterations. This pragma can be used for fine-grained and coarse-grained parallelization. There are also the pragmas ACCEL pipeline <II=y> for pipelining, and ACCEL tile <factor=z> for strip-mining a loop by z, enabling Merlin to insert other pragmas such as data caching in a loop with smaller trip count, matching the on-chip resources available and reducing off-chip communications. Merlin’s ACCEL cache <array=a> directive transfers all required elements of array $a$ from off-chip to on-chip to perform computations within the specified sub-region.
@@ 2.1. DSE for Pragma Insertion
Tools such as AutoDSE (Sohrabizadeh et al., 2021) and GNN-DSE (Sohrabizadeh et al., 2022) rely on Merlin and perform DSE to find which pragmas to insert and where.
Although Merlin further raises the level of abstraction and easiness to implement complex transformations such as data caching and various types of hardware replication and hence reduces the space of configurations, the space of pragmas and their parameter values is still large: as shown later in Sec. 7, this space quickly reaches billions of feasible designs even for kernels containing only a handful of loops and statements.
Furthermore, to avoid accuracy limitations of analytical performance models, DSE approaches exploring such space typically rely on partially generating the design, at least via HLS, and utilizing the QoR report as the performance estimate for a design. HLS time for highly optimized designs combining various Merlin pragmas (e.g., parallelism and caching) can quickly reach several hours per design, making the search process particularly time-consuming. One may restrict the space of pragmas considered, and especially their parameter range, to reduce search time (Sohrabizadeh et al., 2021). However this comes at the expense of the potentially improved performance that can be reached by considering the complete parameter space.
Yet such DSE remains the state-of-practice approach to deliver high-quality designs, we illustrate its performance merits with three sample benchmarks from linear algebra: GEMM, the dense generalized matrix-multiply, 2mm, which computes the matrix product $D=A*B*C$, and Gramschmidt, which computes the Gram-Schmidt process for QR decomposition. Table 1 below displays the performance (in GigaFlop/s, GF/s) of the original program from PolyBench/C using the Medium dataset size and Large for Gramschmidt, without any pragmas, and optimized by Merlin and the final performance achieved by running AutoDSE to insert Merlin pragmas, given a time budget of 20 hours and a timeout of 3 hours per HLS run. DSE time is reported in minutes.
2mm Gemm Gramsch.
GF/s Time GF/s Time GF/s Time
@@
NLP-DSE 117.48 70 105.18 185 2.34 420
Table 1. Result of AutoDSE
This framework utilizes the HLS compilers as a black box used to select configurations that minimize the objective function. The tool is agnostic of the input program shape and if it detects that Merlin did not apply the pragmas as expected it allows the DSE to prune the design after Merlin has generated the HLS-C code.
By analyzing the space explored by the DSE for these three examples, valuable hints can be observed to improve the method. For 2mm, the fastest design found by AutoDSE is mainly optimized for a single loop body. When AutoDSE tries to optimize the second loop body, it favors the unroll factors to the power of two for the innermost loop and goes directly to the outermost loop. This does not improve performance or create configurations which are pruned. The fact that it does not try the other unrolls factors for the innermost loop before optimizing the other pragmas creates a loss of performance. For Gemm Medium and Gramschmidt Large, the DSE makes it possible to find designs with a good QoR. However the DSE wastes much time exploring too large unroll factors, which generates ponderously long synthesis times without giving any result as the HLS timeout is reached. The time spent increasing the unroll factor for certain pragmas without result does not allow the unroll factor of other pragmas to be increased, which results in missed performance.

### v2

@@ Abstract.
High-Level Synthesis enables the rapid prototyping of hardware accelerators, by combining a high-level description of the functional behavior of a kernel with a set of micro-architecture optimizations as inputs. Such optimizations can be described by inserting pragmas for e.g. pipelining and replication of units, or even higher level transformations for HLS such as automatic data caching using the AMD/Xilinx Merlin compiler. Selecting the best combination of pragmas, even within a restricted set, remains particularly challenging and the typical state-of-practice uses design-space exploration to navigate this space. But due to the highly irregular performance distribution of pragma configurations, typical DSE approaches are either extremely time consuming, or operating on a severely restricted search space.
In this work we propose a framework to automatically insert HLS pragmas in regular loop-based programs, supporting pipelining, unit replication (coarse- and fine-grain), and data caching. We develop an analytical performance and resource model as a function of the input program properties and pragmas inserted, using non-linear constraints and objectives. We prove this model provides a lower bound on the actual performance after HLS. We then encode this model as a Non-Linear Program, by making the pragma configuration unknowns of the system, which is computed optimally by solving this NLP. This approach can also be used during DSE, to quickly prune points with a (possibly partial) pragma configuration, driven by lower bounds on achievable latency. We extensively evaluate our end-to-end, fully implemented system, showing it can effectively manipulate spaces of billions of designs in seconds to minutes for the kernels evaluated.
@@
copyright: noneconference: ; ;
@@ 1. Introduction
High-level synthesis (HLS) (Cong et al., 2011; Zhang et al., 2008) compilers (Xilinx, 2023b; Intel, 2023; Siemens, 2023; Microchip, 2023) and source-to-source compiler for HLS (Xilinx, 2023a; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020) can reduce development time while delivering a good performance for the designs. However, achieving a satisfactory Quality of Results (QoR) often requires design-space exploration (DSE). This is because the design space, including which pragmas to insert and where, can not only contain millions of points, but typically does not present characteristics suitable for fast analytical exploration, such as convexity and regularity. Although the existing DSE methods (Sohrabizadeh et al., 2021, 2022; Wu et al., 2023) can find designs with a good QoR, it comes at a high computation cost: for example, hundreds of designs may be concretely instantiated using HLS to compute its estimated QoR during exploration (Sohrabizadeh et al., 2021).
@@
Our main objective in this work is to provide a system to automatically insert a set of hardware pragmas for HLS, that delivers a good QoR and yet significantly reduces the search time needed to obtain the final design. To address this challenge, we propose NLP-DSE , a framework built on top of the AMD/Xilinx Merlin source-to-source compiler (Xilinx, 2023a). This framework automatically inserts pragmas for unrolling/parallelization, pipelining, tiling and data caching for affine programs (Lattner et al., 2021), prior to HLS. These Merlin pragmas can also be inserted using a DSE approach, such as AutoDSE (Sohrabizadeh et al., 2021) or HARP (Sohrabizadeh et al., 2023a), which we use as reference for our evaluations. However, in contrast to AutoDSE (Sohrabizadeh et al., 2021) and HARP (Sohrabizadeh et al., 2023a), we specifically restrict the class of programs we manipulate to affine programs that are regular loop-based computations. In turn, it enables us to develop a hybrid analytical approach to drive the search, combined with a lightweight DSE to reduce the number of designs actually explored. NLP-DSE preserves, and often even improves, the final QoR of designs produced.
@@
To this end, we create a novel Non-Linear Programming approach to automatically insert pragmas in an existing program. We develop an analytical model combining latency and resources, targeting regular loop-based kernels (PolyBench, [n. d.]), that is parameterized by the pragma configuration. We can then designate the pragma configuration as the unknowns of this model, solving it by NLP to obtain the set of pragmas that minimizes latency. An important design principle of our approach is to ensure that the latency computed is a lower bound on the achievable latency for a given pragma configuration. This enables efficient pruning during DSE: any design predicted to have a latency lower bound higher than the best latency obtained through exploration so far is necessarily slower and does not need exploration. To overcome the fact that optimizing compilers, and the overall HLS toolchain underneath, may not apply optimizations as expected (e.g., due to insufficient resources, or limitations of the compiler’s implementation), we develop a lightweight NLP-based DSE approach, exploring parts of the design space with different types and amounts of hardware parallelism and array partitioning factors. We make the following contributions:
@@
  • We present an analytical performance and resource model specifically built for AMD/Xilinx Vitis and Merlin compilers, which is amenable to optimization via non-linear programming.
@@
  • We prove our model is a lower bound on the final latency of the design, under reasonable hypothesis. This enables fast pruning of the design space: it ensures designs which have a higher latency lower bound than the best design found so far can be safely pruned from the search.
@@
  • We develop an NLP-based DSE approach exploiting this model, targeting regular loop-based kernels, which can significantly outperform DSE-based search approaches such as AutoDSE, delivering equal or better QoR in a significantly less search time.
@@
  • We implement NLP-DSE it as an end-to-end, fully automated system and use it to conduct extensive evaluation on 47 benchmarks including kernels from linear algebra, image processing, physics simulation, graph analytics, datamining, etc. (PolyBench, [n. d.]). Our results show the ability of our approach to find in most cases a better QoR than AutoDSE, in significantly less time. Furthermore, in most instances, our approach outperforms HARP in terms of QoR within a comparable timeframe.
@@
The paper is organized as follows. Section 2 motivates our approach and solution proposed. Section 3 presents our analytical performance and resource model. Section 4 delves into proving it is a lower bound on the final QoR. In Section 5, we introduce a non-linear formulation based on this model to automatically find pragma configurations by NLP optimization. Section 6 presents our lightweight DSE approach. Finally, sections 7 and 10 are devoted to evaluating our method validating the effectiveness of our approach and presenting related work, before concluding.
@@ 2. Background and Motivation
2.1. Pragma-based Optimizations for HLS
This work targets the automatic optimization of FPGA designs using HLS (Cong et al., 2011; Zhang et al., 2008), in particular when using compilers to automatically generate HLS-friendly optimized programs such as the AMD/Xilinx Merlin Compiler (Cong et al., 2016b, a; Xilinx, 2023a). Falcon Computing Solutions developed this source-to-source automation tool for FPGAs, which was acquired by Xilinx in late 2020 and is now open source.
HLS has made FPGA usage more accessible, and many projects are looking to further democratize this field by automating optimizations (Sohrabizadeh et al., 2021; Lai et al., 2019; Huang et al., 2021; Ye et al., 2022; Ferretti et al., 2018). Merlin was developed to improve the performance, reduce the development time of HLS-based designs and simplify the search space. To achieve this goal, it automatically generates data transfers between off-chip and on-chip memory and inserts important Vitis pragmas, such as for array partitioning. Moreover, Merlin does specific code transformation for the targeting hardware in adequacy with the hardware directive describe by the user. Additionally, it enables parallel coarse-grained processing elements by encapsulating inner loops as functions.
Merlin includes hardware directive pragmas such as:
  • ACCEL parallel <factor=x>, which creates x parallel instances, and Merlin restructures the code accordingly if the loop nest has more than x iterations. This pragma can be used for fine-grained and coarse-grained parallelization. Merlin will insert the array partitioning corresponding to the parallelization and optimize memory coalescing accordingly for the memory transfer;
  • ACCEL pipeline <II=y> for pipelining;
  • ACCEL tile <factor=z> for strip-mining a loop by z, enabling Merlin to insert other pragmas such as data caching in a loop with smaller trip count, matching the on-chip resources available and reducing off-chip communications;
  • ACCEL cache <array=a> which transfers all required elements of array $a$ from off-chip to on-chip to perform computations within the specified sub-region. If the user does not specify this pragma, it can be applied automatically by Merlin.
In this work we target the automatic generation of pragmas for the Merlin toolchain, to enable the seamless deployment of optimizations such as array partitioning, off-chip data transfers using bursts, coarse-grain and fine-grain replication, etc. These pragma-directed optimizations are implemented by Merlin on loop-based programs, combining source code transformations and the automatic insertion of Vitis pragmas to drive the HLS process.
We note our approach is not restricted to Merlin, nor a particular version of a toolchain: by adjusting the parameters of the performance model, such as operation latency, resource usage per operation(s), etc. one can easily target other toolchains than the one we evaluate here.
@@ 2.2. DSE for Pragma Insertion
Design-space exploration techniques typically trade-off coverage for speed (Sohrabizadeh et al., 2021; Zuo et al., 2015b). That is, it may impose restrictions on the input programs supported, on the pragmas/transformations considered, etc. in order to accelerate the search (Zuo et al., 2015b; Pouchet et al., 2013a). To overcome the difficulty of providing accurate performance models for arbitrary programs, HLS may be invoked to obtain a QoR estimate, without imposing any restriction on the input programs and transformations used. However HLS time for highly optimized designs combining various Merlin pragmas (e.g., parallelism and caching) can quickly reach several hours per design, making the search process particularly time-consuming. One may restrict the space of pragmas considered, and especially their parameter range, to reduce search time. General-purpose DSE approaches such as AutoDSE (Sohrabizadeh et al., 2021) are by design agnostic to the input program features and search space explored, preserving generality. But as we demonstrate in this paper, it also misses opportunities for search acceleration that can be provided by careful static analysis, leading to missed performance opportunities.
In this work, we target the specialization of the DSE process to affine programs, that are programs with a statically analyzable control-flow and dataflow. By restricting the class of programs supported to affine programs, we can deploy exact loop and data dependence analysis (Feautrier, 1992). More importantly, as shown in Section 5, for this class of programs we can model accurately enough the behavior of a design in terms of latency and resource usage by using non-linear programming, significantly accelerating the DSE time for such programs by avoiding the need for actual HLS estimation in numerous cases.
We illustrate the performance merits and limitations of such general-purpose DSE (Sohrabizadeh et al., 2021) on three important loop-based linear algebra benchmarks. GEMM, the classical dense general matrix-multiply, and 2mm shown in Listing 1 which computes the product of three matrices $D=alpha*A*B*C+beta*D$. Both are key computations in e.g., inference of transformers (Devlin et al., 2018). Gramschmidt computes QR decomposition using the Gram-Schmidt process. In later Section 7 we evaluate these benchmarks using various problem sizes, ranging from kBs to MBs of footprints for the matrices to demonstrate the robustness of our approach to varying and large problem sizes. Below we use matrices of about 300kB each.
1 Loop0: for (i1 = 0; i1 < 180; i1++)
2 Loop1: for (j1 = 0; j1 < 190; j1++) {
3 S0: tmp[i1][j1] = 0.0;
4 Loop2: for (k1 = 0; k1 < 210; ++k1)
5 S1: tmp[i1][j1] += alpha * A[i1][k1] * B[k1][j1];
6 }
7 Loop3: for (i2 = 0; i2 < 180; i2++)
8 Loop4: for (j2 = 0; j2 < 220; j2++) {
9 S2: D[i2][j2] *= beta;
10 Loop5: for (k2 = 0; k2 < 190; ++k2)
11 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
12 }
13}
Listing 1: 2mm code: $D=\alpha\times A\times B\times C+\beta\times D$
The search spaces considered here quickly reach billions of feasible designs, even for kernels containing only a handful of loops and statements. Considering 2mm, each loop can have a pragma tile and parallel, all with factors that are divisors of the loop trip count and a pragma pipeline. We obtain a space of $1.37\times 10^{10}$ valid designs. This represents 432 years if assuming one design takes a single second to evaluate. Obviously, only a minimal fraction of these spaces is actually explored, making it essential to adequately select the order in which designs are explored.
Table 1 displays the performance (in GigaFlop/s, GF/s) of the original programs, from PolyBench/C (PolyBench, [n. d.]), when fed to Merlin as-is. The best design found by AutoDSE, given a time budget of 20 hours per benchmark and a timeout of 3 hours per HLS run, is also reported. AutoDSE uses a bottleneck-driven search approach, which targets the improvement of the code section with the lowest throughput (Sohrabizadeh et al., 2021). It unambiguously achieves particularly solid improvements over a naive design without pragmas. However, we show in Table 3 below that a carefully built DSE technique, exploiting the regularity of affine programs and leveraging non-linear programming, can provide order(s) of magnitude higher performance for these exact benchmarks, all while using significantly less search time.
2mm (footprint: 773kB) Gemm (footprint: 579kB) Gramsch. (footprint: 15MB)
GF/s Time (min) GF/s Time (min) GF/s Time (min)
@@ 11 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
Improvement 4.1x 984x 6.8x
Table 1. Comparison of throughput (GF/s) between the AutoDSE framework and the source-to-source compiler Merlin without pragma insertion for the kernels 2mm, Gemm and Gramschmidt
2.3. Limitations of General-Purpose DSE
By analyzing the space explored by the DSE for these three examples, valuable hints can be observed which drive the design of NLP-DSE .
Exploration of the space
AutoDSE (Sohrabizadeh et al., 2021) utilizes the HLS compilers as a black box, in order to select the configurations that minimize the objective function. The tools are agnostic of the input program shape and if AutoDSE detects that Merlin did not apply the pragmas as expected it allows the DSE to prune the design after Merlin has generated the HLS-C code. These frameworks use incremental DSEs, i.e., having no information on the characteristics of the program, they explore the space by increasing the parallelism in order to respond to a problem, e.g., a bottleneck for Auto-DSE.
Table 2 shows the number of valid designs in each space and the number of synthesized, pruned, timeout designs for each kernel. As we can see with a timeout of 20 hours for the DSE and a timeout of 3 hours for each synthesis, the DSE only allows a tiny part of the space to be explored.
2mm Gemm Gramsch.
Nb. valid designs (Space) $1.37\times 10^{10}$ $2.30\times 10^{6}$ $1.22\times 10^{8}$
Nb. design Synthetized (AutoDSE) 15 25 15
Nb. design pruned (AutoDSE) 49 34 239
Nb. design timeout (AutoDSE) 37 27 11
Nb. Design explored (AutoDSE) 101 86 265
Table 2. Investigation into the design space and exploration extent concerning synthesized designs, pruned designs, and designs reaching timeout by the AutoDSE framework for the kernels 2mm, Gemm, and Gramschmidt.
Over Parallelization
AutoDSE is an incremental method: in order to speed up the search AutoDSE will seek to pipeline certain loops which leads to an unrolling of the innermost loops. Without knowledge of the code, trip counts and resources used this leads to over-use of parallelism, leading to timeouts and/or over-use of resources. For 2mm, it attempts to pipeline the outermost loops, leading to the above issues.
Parallelism imbalance
Botlleneck analysis will make it possible to select which part of the code to optimize as a priority (Sohrabizadeh et al., 2021). However, this priority does not take into account parallelism (i.e., hardware resources) that shall be deployed for other parts of the code. This creates code with regions that are extremely/fully parallelized, and others without any parallelism.
For 2mm, the fastest design found by AutoDSE mainly optimizes one loop body. When AutoDSE tries to optimize the second loop body, it favors the unroll factors to the power of two for the innermost loop and goes directly to the outermost loop. This does not improve performance or create configurations which are pruned. The fact that it does not try the other unrolls factors for the innermost loop before optimizing the other pragmas creates a loss of performance. For Gemm and Gramschmidt, the DSE finds designs with a good QoR. However the DSE wastes much time exploring too large unroll factors, which generates ponderously long synthesis times without giving any result as the HLS timeout is reached. The time spent increasing the unroll factor for certain pragmas without result does not allow the unroll factor of other pragmas to be increased, which results in missed performance.